# ElaipBench: A Benchmark for Expert-Level Artificial Intelligence Paper Understanding

### v1

@@ Abstract
While large language models (LLMs) excel at many domain-specific tasks, their ability to deeply comprehend and reason about full-length academic papers remains underexplored. Existing benchmarks often fall short of capturing such depth, either due to surface-level question design or unreliable evaluation metrics. To address this gap, we introduce ElaipBench, a benchmark curated by domain experts to evaluate LLMs’ comprehension of artificial intelligence (AI) research papers. Developed through an incentive-driven, adversarial annotation process, ElaipBench features 403 multiple-choice questions from 137 papers. It spans three difficulty levels and emphasizes non-trivial reasoning rather than shallow retrieval. Our experiments show that the best-performing LLM achieves an accuracy of only 39.95%, far below human performance. Moreover, we observe that frontier LLMs equipped with a thinking mode or a retrieval-augmented generation (RAG) system fail to improve final results—even harming accuracy due to overthinking or noisy retrieval. These findings underscore the significant gap between current LLM capabilities and genuine comprehension of academic papers. [^1] The benchmark is publicly available at https://huggingface.co/datasets/KangKang625/ELAIPBench. Building on this work, we organized the CCKS-2025 Complex Question Answering Competition for AI Papers using a subset of ElaipBench. The final rankings are accessible at https://tianchi.aliyun.com/competition/entrance/532359/rankingList.
@@ 1 Introduction
Unfortunately, existing domain-specific benchmarks fail to adequately assess the deep comprehension of LLMs for academic papers. Some benchmarks focus on content extraction tasks where answers are explicitly stated in the papers Lála et al. (2023); Auer et al. (2023); Laurent et al. (2024). These answers can be easily retrieved by LLMs due to their strong recall abilities. Others employ open-ended questions requiring the summarization of phenomena or methods described in the papers Saikh et al. (2022); Lee et al. (2023); Wan et al. (2024); Wellawatte et al. (2025); however, their evaluation metrics (e.g., ROUGE) fail to accurately and reliably measure fine-grained details or key technical knowledge. Crucially, existing approaches relying on human annotation struggle with quality control—although some incorporate question validation mechanisms Welbl et al. (2017); Bai et al. (2024), crowd-sourced annotation schemes may still produce superficial or low-quality questions Kittur et al. (2008); Snow et al. (2008).
@@
To ensure the difficulty and high quality of ElaipBench, we employ a competitive incentive annotation process to create challenging questions. We recruit 20 human annotators, each holding at least a master’s degree in computer science and having prior experience in publishing academic papers. Each annotator is assigned one of three roles: Question Writer, Evidence Verifier, or Answer Verifier.The Question Writer is responsible for creating data instances. Each instance comprises an English academic paper, a set of questions with corresponding answer choices, the correct answer, and supporting evidence excerpts from the paper. Subsequently, as a preliminary difficulty filter, we use three LLMs for automated review. If any model answers the question correctly, it is discarded as too easy. The filtered questions are assigned to Evidence Verifiers, who answer them using the paper and evidence excerpts. A question is retained only if the Evidence Verifier derives the correct answer using the provided evidence. Finally, the Answer Verifier attempts to answer the questions based solely on the paper, question, and options. Throughout the process, annotators receive a base salary and compete for performance bonuses based on response time and accuracy. This mechanism categorizes questions into three difficulty levels and motivates annotators to create challenging yet answerable questions. Figure 1 presents an example question from ElaipBench. Its options not only incorporate explicit content from the paper but also require LLMs to perform reasoning based on this content combined with scientific common knowledge. Overall, our contributions are summarized as follows:
@@
  • We propose a competitive incentive annotation mechanism that uses performance-linked bonuses to generate high-quality questions. Comprehensive evaluations show the benchmark’s difficulty: human experts achieve only 48.14% accuracy within 20 minutes, while the best-performing LLMs reach just 39.95%.
@@ 2.1 Scientific Comprehension Benchmarks
To advance the evaluation of LLMs in the domain of scientific comprehension, a variety of benchmark datasets have emerged, with increasing emphasis on task complexity and domain specificity. These benchmarks can be broadly categorized into two groups. The first group either does not require contextual information or relies on only brief text passages. Sorted by publication date, these benchmarks include emrQA Pampari et al. (2018), PubMedQA Jin et al. (2019), ScienceQA Saikh et al. (2022), SciQA Auer et al. (2023), SciBench Wang et al. (2023), ChemLit-QA Wellawatte et al. (2025), and HLE Phan et al. (2025). They primarily assess models’ intrinsic knowledge within a specific domain while neglecting the critical ability to comprehend contextual information—an essential skill in academic research or paper review. The second group incorporates full academic papers as context Lála et al. (2023); Skarlinski et al. (2024); Bai et al. (2024); however, the associated questions tend to be superficial, enabling LLMs to locate answers through direct text matching rather than requiring synthesis, inference, or critical engagement with the academic paper context.
2.2 Academic Benchmark Annotation Methods
Current academic question answering benchmarks are primarily generated using two annotation approaches. The first relies on crowdsourced human annotation, where data labeling is conducted through pipeline-based workflows and quality control is typically managed by a review group Lála et al. (2023); Bai et al. (2024); Skarlinski et al. (2024); Asai et al. (2024). This approach, however, heavily depends on the efficiency and expertise of the reviewing team, with limited mechanisms for providing feedback on the review outcomes. The second approach leverages LLMs to automatically generate datasets by equipping them with various specialized tools and prompting frameworks Lee et al. (2023); Wan et al. (2024); Kim et al. (2025); Yu et al. (2025). While scalable, this method is constrained by the generative model’s own comprehension ceiling. The resulting questions, therefore, often lack diversity in reasoning types and fail to achieve sufficient depth in academic problem-solving.
@@ 3 Construction of ElaipBench
In this section, we present the construction process of ElaipBench. First, we define the task overview. Next, we introduce the incentive competition annotation mechanism designed for generating high-quality, challenging data. Finally, we provide statistical information about the benchmark.
@@ 3.1 Task Overview
3.2 Competitive Incentive Annotation Mechanism
As illustrated in Figure 2, our incentive-driven competitive mechanism consists of three key stages: QA pair generation, Evidence Verification, and Answer Verification. These stages involve three distinct roles (Question Writer, Evidence Verifier, and Answer Verifier), staffed with 10, 4, and 6 annotators, respectively. The roles operate independently yet pursue competing objectives within the workflow. Compensation is structured as a combination of base pay and performance-based bonuses. Annotators receive a base payment of 30 CNY per question they generate. During the verification stages, they could earn additional performance bonuses of 30 CNY (Level-1) or 60 CNY (Level-2), awarded based on the quality of their annotations. The total cost for all data collection and annotation was approximately 50,000 CNY.
@@ 3.2.2 Evidence Verification
In this stage, Evidence Verifiers assess whether the provided evidence excerpts are sufficient to support a correct response. This requires: (1) analyzing and reasoning over the evidence, and (2) attempting to answer the question using only this evidence. To incentivize emphQuestion Writers to provide accurate and logically aligned evidence, we implement a competitive mechanism: if the Evidence Verifier arrives at the correct answer using exclusively the provided evidence, both receive Level-1 performance bonuses. Conversely, if verification fails, penalties apply: the Question Writer and the Evidence Verifier lose their bonus allocation. This failure triggers a Revision cycle where the Question Writer must improve either the evidence or the entire QA pair. The revised version must pass both Automatic Review and subsequent Evidence Verification before either party is rewarded. This dual-incentive mechanism simultaneously maintains question difficulty standards and compels Question Writers to provide logically sound evidence supporting the answer. During the annotation process, 67% of the questions initially failed the Automated Review, and 38% failed the initial Evidence Verification.
@@ 3.2.3 Answer Verification
This mechanism establishes a dynamic adversarial relationship between the Question Writer and the Answer Verifier: Question Writers are incentivized to maximize question difficulty for higher compensation, while Answer Verifiers are motivated to provide accurate responses efficiently to secure their bonuses.
@@ 4 Experiments
Human - 48.14 56.82 45.71 100.00 100.0 0.00 45.74 51.85 52.20
Table 2: Accuracy of LLMs on the ElaipBench. Bold indicates the best result among LLMs. The Human results were given by the Answer Verifier during the annotation process.
@@ 4.2 Results for LLMs
We employ three primary paradigms: a base model paradigm, a chain-of-thought (CoT) paradigm, and a LRM paradigm. In the base model paradigm, models are instructed to output only the final answer option. In the CoT paradigm, we add the prompt “Let’s think step by step.” to elicit stepwise reasoning before the final answer. In the LRM paradigm, the model is allowed to produce an unstructured reasoning chain before the answer.
@@ 4.3 Reasoning Error Analysis
Figure 4 illustrates the relationship between reasoning length and accuracy across different LLMs. Although QwQ-32B generates the longest reasoning chains, its accuracy does not improve with increased length—in fact, it achieves the lowest performance among the models. DeepSeek-R1 attains the highest accuracy among the three LRMs while maintaining a relatively short average reasoning chain, demonstrating superior reasoning efficiency. Nevertheless, its performance remains significantly lower than that of Qwen3-32B+CoT, suggesting that longer reasoning chains do not necessarily lead to better performance and may even be detrimental. We observe that self-verification behaviors occupy a substantial portion—approximately 50%—of the LRM’s reasoning chain. These behaviors typically initiate with phrases such as Wait or Alternatively, and are followed by re-statements of the original question and subsequent restarts of the reasoning chain. Such repetitive patterns significantly lengthen the reasoning trajectory, introducing ineffective verification steps. Representative cases are provided in Appendix A.5.
@@ Appendix A Appendix
A.1 Annotation Guidelines
@@
A.2 Annotation Details
A.2.1 Annotator Statistics
@@ A.3 Benchmark Statistics
Table 3 shows the statistics of the papers. Tables 4 through 9 present representative SA-MCQ and MA-MCQ examples from the three disciplines: ML, CV, and NLP.
@@ D. OPT-13B
A.4 Evaluation Metric and LLM Settings
@@
A.5 Error Analysis
@@
Figure 19: An example of Harmful Verifications from QwQ-32B, where the LLM first identifies the correct answer but then incorrectly rejects it during reasoning, ultimately selecting a wrong alternative.
A.6 RAG Settings
A.6.1 Hyper-parameter Setting

### v2

@@ Abstract
While large language models (LLMs) excel at many domain-specific tasks, their ability to deeply comprehend and reason about full-length academic papers remains underexplored. Existing benchmarks often fall short of capturing such depth, either due to surface-level question design or unreliable evaluation metrics. To address this gap, we introduce ElaipBench, a benchmark curated by domain experts to evaluate LLMs’ comprehension of artificial intelligence (AI) research papers. Developed through a game-theoretic, adversarial annotation process, ElaipBench features 403 multiple-choice questions from 137 papers. It spans three difficulty levels and emphasizes non-trivial reasoning rather than shallow retrieval. Our experiments show that the best-performing LLM achieves an accuracy of only 39.95%, far below human performance. Moreover, we observe that frontier LLMs equipped with a thinking mode or a retrieval-augmented generation (RAG) system fail to improve final results—even harming accuracy due to overthinking or noisy retrieval. These findings underscore the significant gap between current LLM capabilities and genuine comprehension of academic papers.  [^1] The benchmark is publicly available at https://huggingface.co/datasets/KangKang625/ELAIPBench. Building on this work, we organized the CCKS-2025 Complex Question Answering Competition for AI Papers using a subset of ElaipBench. The final rankings are accessible at https://tianchi.aliyun.com/competition/entrance/532359/rankingList.
@@ 1 Introduction
Unfortunately, existing domain-specific benchmarks fail to adequately assess the deep comprehension of LLMs for academic papers. Some benchmarks focus on content extraction tasks where answers are explicitly stated in the papers Lála et al. (2023); Auer et al. (2023); Laurent et al. (2024). These answers can be easily retrieved by LLMs due to their strong recall abilities. Others employ open-ended questions requiring the summarization of phenomena or methods described in the papers Saikh et al. (2022); Lee et al. (2023); Wan et al. (2024); Gui et al. (2025); however, their evaluation metrics (e.g., ROUGE) fail to accurately and reliably measure fine-grained details or key technical knowledge. Crucially, existing approaches relying on human annotation struggle with quality control—although some incorporate question validation mechanisms Welbl et al. (2017); Bai et al. (2024), crowd-sourced annotation schemes may still produce superficial questions Kittur et al. (2008).
@@
To ensure the difficulty and high quality of ElaipBench, we employ a game-theoretic competitive annotation protocol designed to yield challenging questions. We recruit 20 human annotators, each holding at least a master’s degree in computer science and having prior experience in publishing academic papers. Each annotator is assigned one of three roles: Question Writer, Evidence Verifier, or Answer Verifier.The Question Writer is responsible for creating data instances. Each instance comprises an English academic paper, a set of questions with corresponding answer choices, the correct answer, and supporting evidence excerpts from the paper. Subsequently, as a preliminary difficulty filter, we use three LLMs for automated review. If any model answers the question correctly, it is discarded as too easy. The filtered questions are assigned to Evidence Verifiers, who answer them using the paper and evidence excerpts. A question is retained only if the Evidence Verifier derives the correct answer using the provided evidence. Finally, the Answer Verifier attempts to answer the questions based solely on the paper, question, and options. Throughout the process, annotators receive a base salary and compete for performance bonuses based on response time and accuracy. By adopting this game-theoretic incentive design, we establish careful annotation as the Nash equilibrium for all annotators, thereby yielding challenging yet answerable questions. Figure 1 presents an example question from ElaipBench. Its options not only incorporate explicit content from the paper but also require LLMs to perform reasoning based on this content combined with scientific common knowledge. Overall, our contributions are summarized as follows:
@@
  • We propose a game-theoretic annotation mechanism that uses performance-linked bonuses to generate high-quality questions. Comprehensive evaluations show the benchmark’s difficulty: human experts achieve only 48.14% accuracy within 20 minutes, while the best-performing LLMs reach just 39.95%.
@@ 2.1 Scientific Comprehension Benchmarks
To advance the evaluation of LLMs in the domain of scientific comprehension, a variety of benchmark datasets have emerged, with increasing emphasis on task complexity and domain specificity. These benchmarks can be broadly categorized into two groups. The first group either does not require contextual information or relies on only brief text passages. Sorted by publication date, these benchmarks include emrQA Pampari et al. (2018), PubMedQA Jin et al. (2019), ScienceQA Saikh et al. (2022), SciQA Auer et al. (2023), SciBench Wang et al. (2023), ChemLit-QA Wellawatte et al. (2025), and HLE Phan et al. (2025). They primarily assess models’ intrinsic knowledge within a specific domain while neglecting the critical ability to comprehend contextual information—an essential skill in academic research or paper review. The second group incorporates full academic papers as context Lála et al. (2023); Skarlinski et al. (2024); Bai et al. (2024); however, the associated questions tend to be superficial, enabling LLMs to locate answers through direct text matching rather than requiring synthesis, inference, or critical engagement with the academic paper context. For more detailed comparisons of existing benchmarks, please refer to Appendix A.1.
2.2 Annotation Methods
Current academic question answering benchmarks are primarily generated using two annotation approaches. The first relies on crowdsourced human annotation, where data labeling is conducted through pipeline-based workflows and quality control is typically managed by a review group Lála et al. (2023); Bai et al. (2024); Skarlinski et al. (2024); Asai et al. (2024); Gui et al. (2025). This approach, however, heavily depends on the efficiency and expertise of the reviewing team, with limited mechanisms for providing feedback on the review outcomes. The second approach leverages LLMs to automatically generate datasets by equipping them with various specialized tools and prompting frameworks Lee et al. (2023); Wan et al. (2024); Kim et al. (2025); Yu et al. (2025). While scalable, this method is constrained by the inherent comprehension limits of the models, frequently resulting in insufficient diversity and a lack of depth required for resolving complex academic questions. To the best of our knowledge, we are the first to integrate game-theoretic principles into the annotation process for academic QA benchmarks.
@@ 3 Construction of ElaipBench
In this section, we present the construction process of ElaipBench. First, we define the task. Next, we introduce a game-theoretic annotation mechanism designed to ensure challenging data. Finally, we provide statistical information.
@@ 3.1 Task Overview
3.2 Game-Theoretic Annotation Mechanism
As illustrated in Figure 2, our annotation mechanism consists of three key stages: QA Pair Generation, Evidence Verification, and Answer Verification. These stages involve three distinct roles (Question Writer, Evidence Verifier, and Answer Verifier), staffed with 10, 4, and 6 annotators, respectively. The roles operate independently yet pursue competing objectives within the workflow. Compensation is structured as a combination of base pay and performance-based bonuses. Annotators receive a base payment of 30 CNY per question they generate. During the verification stages, they can earn additional performance bonuses of 30 CNY (Level-1) or 60 CNY (Level-2), awarded based on the quality of their annotations. The total cost for all data collection and annotation was over 50,000 CNY (~7,000 USD).
@@ 3.2.2 Evidence Verification
In this stage, Evidence Verifiers assess whether the provided evidence excerpts are sufficient to support a correct choice. This requires: (1) analyzing and reasoning over the evidence, and (2) attempting to answer the question using only this evidence. To incentivize Question Writers to provide accurate and logically aligned evidence, we implement a competitive mechanism: if the Evidence Verifier arrives at the correct answer using exclusively the provided evidence, both receive Level-1 performance bonuses. Conversely, if verification fails, penalties apply: the Question Writer and the Evidence Verifier lose their bonus allocation. This failure triggers a Revision cycle where the Question Writer must improve either the evidence or the entire QA pair. The revised version must pass both Automatic Review and subsequent Evidence Verification before either party is rewarded. This filtering mechanism ensures that questions clearing this stage are fully answerable by human experts given the evidence excerpts. Crucially, this game-theoretic protocol compels Question Writers to simultaneously elevate the difficulty standards of their questions while ensuring the provision of logically sound evidence to support their answers.
@@ 3.2.3 Answer Verification
In summary, the proposed mechanism induces a dynamic adversarial relationship among the three annotator groups and establishes careful annotation as a Nash equilibrium. Specifically, any deviation from diligent annotation is identified through cross-validation against the others’ annotations, consequently resulting in reduced compensation.
@@ 4 Experiments
Human - 48.14 56.82 45.71 100.00 100.00 0.00/100.00* 45.74 51.85 52.20
Table 2: Accuracy of LLMs on the ElaipBench. Bold indicates the best result among LLMs. The Human results are provided by the Answer Verifier, and values marked with * indicate the accuracy of the Evidence Verifier.
@@ 4.2 Results for LLMs
We employ three primary paradigms: a base model paradigm, a chain-of-thought (CoT) paradigm, and an LRM paradigm. In the base model paradigm, models are instructed to output only the final answer option. In the CoT paradigm, we add the prompt “Let’s think step by step.” to elicit stepwise reasoning before the final answer. In the LRM paradigm, the model is allowed to produce an unstructured reasoning chain before the answer.
@@ 4.3 Reasoning Error Analysis
Figure 4 shows the relationship between reasoning length and accuracy across different LRMs. A consistent trend emerges: as the output length increases from shorter to longer ranges, accuracy does not improve for any model; instead, it steadily declines. This degradation is most pronounced for Claude-3.7-Sonnet-thinking. These results at least indicate that longer outputs are not a reliable indicator of more thorough reasoning; rather, they often reflect a tendency to generate additional explanations, branching hypotheses, or redundant steps when the model is uncertain. We further observe that self-verification behaviors account for a substantial fraction of LRM reasoning traces (approximately 50%). Such segments often begin with phrases such as wait or Alternatively, followed by a restatement of the original question and a restart of the reasoning process. This repetitive pattern substantially lengthens the reasoning path while introducing largely ineffective verification steps. Representative cases are provided in Appendix A.6, together with statistics summarizing the average reasoning length across models.
@@ Appendix A Appendix
A.1 Academic Paper Understanding Dataset Comparison
As shown in Table 3, we summarize key dataset characteristics along multiple dimensions, including whether the benchmark is constructed from full-length papers, the number of instances, task format, the presence of objective evaluation metrics, and whether the data are human-annotated.
Datasets Full-Length Paper #Data Task Objective Evaluation Metric Human Annotation
ArxivBench Li et al. (2025) 6,500 Open QA
SciQAG Wan et al. (2024) 188,042 Open QA
ChemLit-QA Wellawatte et al. (2025) 1,000 Open QA
ScienceQA Saikh et al. (2022) 100k Open QA
LITQA Lála et al. (2023) 50 MCQ
PaperBench Starace et al. (2025) 20 Code Generation
AcadReason Gui et al. (2025) 50 Open QA
ELAIPBench 403 MCQ
Table 3: Academic paper understanding dataset.
Among the listed benchmarks, only a small subset—PaperBench Starace et al. (2025), AcadReason Gui et al. (2025), and our proposed ELAIPBench—is built on complete scientific papers. This distinction is crucial: understanding an entire paper requires sustained comprehension of complex methodology, nuanced argumentation, and long-range dependencies, which cannot be adequately assessed using only abstracts, snippets, or short prompts. By contrast, ArxivBench Li et al. (2025), SciQAG Wan et al. (2024), ChemLit-QA Wellawatte et al. (2025), and ScienceQA Saikh et al. (2022) rely on short or automatically generated inputs that may measure retrieval or surface-level knowledge, but are insufficient for evaluating deep understanding over full academic texts.
Dataset scale further differentiates existing resources. While some benchmarks (e.g., SciQAG) contain more than 180,000 instances, these are automatically generated and often lack the authenticity and depth required to assess expert-level reasoning. In contrast, human-annotated datasets are typically much smaller due to high cost and domain expertise requirements. For example, LITQA Lála et al. (2023) contains only 50 MCQs and does not provide full-paper context; PaperBench includes only 20 instances and primarily targets code generation rather than comprehension. Although PaperBench places substantial demands on planning, engineering, and coding ability, these skills are largely orthogonal to the central goal of evaluating scientific understanding. AcadReason is based on full papers but likewise includes only 50 questions, limiting its statistical power for robust model evaluation.
Task design and evaluability also vary substantially across benchmarks. Many datasets, including ArxivBench, SciQAG, and AcadReason, adopt open QA settings, making objective and reproducible evaluation difficult. Accordingly, we approach such paradigms with caution when aiming for rigorous benchmarking.
Finally, human annotation remains essential for producing high-quality, domain-specific questions grounded in genuine scientific reasoning. Although automatic generation can scale efficiently, it often falls short in the subtlety and cognitive rigor needed for expert assessment. Our annotation protocol further strengthens reliability via a game-theoretic mechanism that incentivizes careful reading and verification, thereby excluding superficial engagement that can arise in pipeline-style labeling.
In summary, ELAIPBench uniquely synthesizes four key attributes: (1) the utilization of full-text papers, (2) large-scale human annotation, (3) objective evaluation via MCQs, and (4) a focus on deep comprehension rather than citation retrieval or code generation. These properties make ELAIPBench a rigorous and reliable tool for evaluating advanced LLMs in realistic academic comprehension scenarios.
A.2 Annotation Guidelines
@@
A.3 Annotation Details
A.3.1 Annotator Statistics
@@ A.4 Benchmark Statistics
Table 4 shows the statistics of the papers. Tables 5 through 10 present representative SA-MCQ and MA-MCQ examples from the three disciplines: ML, CV, and NLP.
@@ D. OPT-13B
A.5 Evaluation Metric and LLM Settings
@@
Figure 19: An example of Harmful Verification from QwQ-32B, where the LLM first identifies the correct answer but then incorrectly rejects it during reasoning, ultimately selecting a wrong alternative.
A.6 Error and Efficiency Analysis
@@
Figure 21: Completion Tokens vs.Performance.
Figure 21 illustrates the relationship between average reasoning length and accuracy across different LLMs. Although QwQ-32B generates the longest reasoning chains, its accuracy does not improve with increased length—in fact, it achieves the lowest performance among the models. DeepSeek-R1 attains the highest accuracy among the three LRMs while maintaining a relatively short average reasoning chain, demonstrating superior reasoning efficiency. Nevertheless, its performance remains significantly lower than that of Qwen3-32B+CoT, suggesting that, in many models, longer reasoning chains do not necessarily lead to better performance and may even be detrimental.
A.7 RAG Settings
A.7.1 Hyper-parameter Setting